OpenStudy (olivya999):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (olivya999):

@Kainui @sleepyjess

OpenStudy (olivya999):

@mayankdevnani @Atsie @AaronAndyson @Anna1448 @Ashely14 @Bobo-i-bo @birdy258 @carolina4567

OpenStudy (olivya999):

@Zenmo @4815162342

OpenStudy (4815162342):

4.8

OpenStudy (olivya999):

that's the answer?

OpenStudy (4815162342):

probably

OpenStudy (olivya999):

@Qwertty123 can you help?

OpenStudy (olivya999):

how do you know?

OpenStudy (4815162342):

I used the calculator

OpenStudy (zenmo):

Delete all your replies, I did the equation wrong.

OpenStudy (olivya999):

i just did @Zenmo

OpenStudy (bobo-i-bo):

Hmmm, so the simpliest/brute method would be to: \[7sin(2x)-14sinx+2=-5\] \[14sinxcosx-14sinx=-7\] \[14\sqrt{1-sin^2x}sinx=14sinx-7\] \[14^2(1-sin^2x)sin^2x=(14sinx-7)^2\] Then just expand and simply and see if you can solve the resulting quartic polynomial

Nnesha (nnesha):

I think it is sin squared x

OpenStudy (olivya999):

how do i expand?

Nnesha (nnesha):

It is quadratic equation ,

Nnesha (nnesha):

I'm on phone can't type

Nnesha (nnesha):

Set the equation equal to and then solve the quadratic equation

Nnesha (nnesha):

Equal to 0*

Nnesha (nnesha):

Equal to 0*

Nnesha (nnesha):

7sin^2x-14sinx +2 +5=0

Nnesha (nnesha):

Let sin x=y, 7y^2 - 14y+7=0 , take out the common factor & factor

jhonyy9 (jhonyy9):

@olivya999 here in OS we like learn you how you can solving same exercises not direct answers - ok . ?

OpenStudy (olivya999):

Okay learn me @jhonyy9

jhonyy9 (jhonyy9):

ok. so how have wrote to you above @Nnesha you sign sin x = y and will get one quadratic in y what you need solve it for y do you can solve it please ?

OpenStudy (olivya999):

Y=1 @jhonyy9

jhonyy9 (jhonyy9):

so than this is right you need push sin x = 1 and calculi the value of x